OpenStudy (anonymous):

What is the value of b2 - 4ac for the following equation? 5x2 + 7x = 6 9 81 169

jhonyy9 (jhonyy9):

so than b=7 a=5 and c=-6 can you calculi it now ?

OpenStudy (anonymous):

7^2-(4*5*-6)=169

jhonyy9 (jhonyy9):

so b^2 = 49 -4ac=-4*5*(-6)=-20(-6)=120 120!49=169

jhonyy9 (jhonyy9):

sorry 120+49=169

OpenStudy (anonymous):

I have another one. 2X^2-2x-1=0